OpenStudy (anonymous):

solve. do not forget to check 5 =6 - 1 -- -- -- x x 3

OpenStudy (anonymous):

\[5/x=(18-x)/3x\] 15x=18x-x^2 3x-x^2=0 x(3-x)=0 this will five you: x=0 and x=3 but! x can not be =0 because you can not divide by 0. so, there is only one answer: x=3

jhonyy9 (jhonyy9):

15=18-x --- x=3